# Cross-view Action Recognition Understanding From Exocentric to Egocentric Perspective

### v2

@@ 2 Related Work
Video Action Recognition Many large-scale third-view datasets have been introduced for action recognition tasks, e.g., Kinetics [13, 18, 19], Something-Something V2 [20], Sport1M [21], AVA [22], etc. Many deep learning approaches [2, 3, 15, 23, 24, 25, 26, 27] have been introduced and achieved remarkable achievements. The early deep learning approaches [21, 28] have utilized the 2D Convolutional Neural Networks (CNNs) [29, 30, 31] to extract the deep spatial representation followed by using Recurrent Neural Networks (RNNs) [32] to learn the temporal information from these extracted spatial features. Some later approaches have improved the temporal learning capability by introducing the two-stream networks [24, 33, 34, 35, 36] using both RGB video inputs and optical flows for motion modeling. Later, the 3D CNN-based approaches [37] and their variants [23, 38] have been introduced, i.e., several (2+1)D CNN architectures have been proposed [15, 39, 40, 41]. Meanwhile, other approaches have used pseudo-3D CNNs built based on 2D CNNs [25, 42]. In addition, to better capture the long-range temporal dependencies among video frames, the non-local operation has also been introduced [43]. SlowFast [15] proposes a dual-path network to learn spatiotemporal information at two different temporal rates. X3D [40] progressively expands the networks to search for an optimal network for action recognition.
@@
Vision Transformer [2, 3, 26, 27, 44, 45, 46] has become a dominant backbone in various tasks due to its outstanding performance. The early success of Video Vision Transformer (ViViT) [26] has shown its promising capability in handling spatial-temporal tokens in action recognition. Then, many variants [2, 3, 27, 45, 46, 47, 48] of ViViT have been introduced to improve the accuracy and reduce the computational cost. [49] presented a space-time mixing attention mechanism to reduce the complexity of the self-attention layers. TimeSFormer [45] introduced divided spatial and temporal attention to reduce the computational overhead. Then, it is further improved using the directed attention mechanism [46]. Then, [27] proposed a Multi-scale Vision Transformer (MViT) using multiscale feature hierarchies. Then, MViT-V2 [3] improves the performance of MViT by incorporating decomposed relative positional embeddings and residual pooling connections. Swin Video Transformer [2] has achieved state-of-the-art performance in action recognition by using shifted windows to limit the self-attention computational cost to local windows and also allow learning attention across windows. The recent SVFormer [50] has introduced a temporal warping augmentation to capture the complex temporal variation in videos for semi-supervised action recognition. Meanwhile, MTV [51] presented a Multiview Transformer to model different views of the videos with lateral connections to fuse information across views. Later, M&M Mix[52] further improved MTV by using multimodal inputs. TADA [53] proposed Temporally-Adaptive Convolutions (TAdaConv) to model complex temporal dynamics in videos. Inspired by the success of CLIP in vision-language pretraining [54, 55, 56], several studies have adopted this approach to video-language pretraining [5, 57, 58]. All-in-One [57] presented a unified approach to video-language pretraining by embedding raw video and textual inputs into joint representations with a unified network. EgoVLP [5] introduced Video-Language Pretraining to ego-centric video understanding. LF-VILA [58] presented a Multimodal Temporal Contrastive and Hierarchical Temporal Window Attention to model the long-form videos for video-language pretraining.
@@
Cross-view Video Learning The cross-view learning approaches have been exploited and proposed for several tasks, e.g., geo-localization [80, 81, 82, 83], semantic segmentation [84, 85, 86]. Meanwhile, in video understanding tasks, several prior methods have alleviated the cross-view gap between exocentric and egocentric domains by using domain adaptation [16, 87, 88], learning viewpoint-invariant [17, 89, 90, 91], or learning joint embedding [92, 93, 94, 95, 96]. Other works utilized generative models to synthesize the different viewpoints from a given image/video [97, 98, 99, 100]. However, these methods often require either a pair of data of both first and third views to learn the joint embedding or a share label domain when using domain adaptation.
@@ 4.2 Ablation Studies
Effectiveness of the scale $\alpha$ We study the effectiveness of the linear scale $\alpha$ to the performance of the model. In this experiment, the metrics defined in Eqn. (12) and Eqn. (13) have been adopted to $\mathcal{D}_{x}$ and $\mathcal{D}_{a}$. The value $\alpha$ ranges from 0.0 to 2.0. When $\alpha=0.0$, it is equivalent to ViT simultaneously trained on both third-view and first-view datasets. As shown in Table 1, the mAP performance on the Charades-Ego benchmark is consistently improved when the value of $\alpha$ increases from $0.1$ to $0.75$ and achieves the best performance at the value of $\alpha=0.75$ and the mAP performance is $31.95\%$. Similarly, on the EPIC-Kitchen-55 benchmarks, the Top 1 and Top 5 accuracy is gradually improved w.r.t the increasing of $\alpha$ and reaches the maximum performance when the value of $\alpha$ is $1.50$ in which the Top 1 accuracy on EPIC Verb and EPIC Noun are $73.52\%$ and $68.19\%$. Then, the performance on both benchmarks steadily decreases when the value of $\alpha$ keeps increasing over the optimal point. Indeed, the variation in the video space is typically higher than in the attention maps due to the higher complexity of video data where the video data contains much more information, e.g., objects, humans, and interactions, etc.; meanwhile, the attention maps represent the focus of the models w.r.t model decisions. Thus, if the value of $\alpha$ is small, it could not represent the correct proportion of changes between videos and attention maps. Meanwhile, the higher value of $\alpha$ inclines to exaggerate the model focuses, i.e., attention maps, that results in the performance.
@@
Effectiveness of Transformer Layers This experiment studies the effectiveness of imposing the cross-view loss into attention maps of the Transformer layers. In this experiment, we adopt the optimal setting of the linear scale ($\alpha$) and correlation metrics ($\mathcal{D}_{x}$, $\mathcal{D}_{a}$) in the previous ablation studies. We consider four groups of Transformer layers, each consisting of three consecutive layers, i.e., Layer 1-3, Layer 4-6, Layer 7-9, and Layer 10-12. As experimental results in Table 3, the later Transformer layers of our model play an important role than the initial ones. In particular, when imposing the cross-view loss on only the first three Transformer layers, the performance of Charades-Ego has achieved 25.65% and the Top 1 accuracy of verb and noun predictions in EPIC-Kitchens-55 is 60.47% and 57.85%. Meanwhile, enforcing the cross-view self-attention loss into all attention layers brings better performance and achieves the best performance, i.e., the mAP of 31.95% on Charades-Ego and Top 1 accuracy of 73.52% and 68.19% on EPIC-Kitchens-55. Fig. 4 visualizes the attention maps of the model predictions.
@@
Table 5: Comparisons on EPIC-Kitchen-55.
Method EPIC verbs EPIC nouns
Top 1 Top 5 Top 1 Top 5
ResNet-50 [15] 61.19 87.49 46.18 69.72
MViT-V2 [3] 55.17 89.87 56.59 79.40
Swin-B [2] 56.40 85.84 47.68 71.02
DANN [16] 61.27 87.49 45.93 68.73
Joint-Embed [17] 61.26 87.17 46.55 68.97
Ego-Exo + ResNet-50 [1] 62.83 87.63 48.15 70.28
Ego-Exo + SlowFast [1] 65.97 88.91 49.42 72.35
Ego-Exo* + ResNet-50 [1] 64.26 88.45 48.39 70.68
Ego-Exo* + SlowFast [1] 66.43 89.16 49.79 71.60
CVAR (Ours) 73.52 92.22 68.19 84.93

### v3

@@ 2 Related Work
Video Action Recognition Many large-scale third-view datasets have been introduced for action recognition tasks, e.g., Kinetics [13, 18, 19], Something-Something V2 [20], Sport1M [21], AVA [22], etc. Many deep learning approaches [2, 3, 15, 23, 24, 25, 26, 27, 28] have been introduced and achieved remarkable achievements. The early deep learning approaches [21, 29] have utilized the 2D Convolutional Neural Networks (CNNs) [30, 31, 32] to extract the deep spatial representation followed by using Recurrent Neural Networks (RNNs) [33] to learn the temporal information from these extracted spatial features. Some later approaches have improved the temporal learning capability by introducing the two-stream networks [24, 34, 35, 36, 37] using both RGB video inputs and optical flows for motion modeling. Later, the 3D CNN-based approaches [38] and their variants [23, 39] have been introduced, i.e., several (2+1)D CNN architectures have been proposed [15, 40, 41, 42]. Meanwhile, other approaches have used pseudo-3D CNNs built based on 2D CNNs [25, 43]. In addition, to better capture the long-range temporal dependencies among video frames, the non-local operation has also been introduced [44]. SlowFast [15] proposes a dual-path network to learn spatiotemporal information at two different temporal rates. X3D [41] progressively expands the networks to search for an optimal network for action recognition.
@@
Vision Transformer [2, 3, 26, 27, 45, 46, 28, 47, 48] has become a dominant backbone in various tasks due to its outstanding performance. The early success of Video Vision Transformer (ViViT) [26] has shown its promising capability in handling spatial-temporal tokens in action recognition. Then, many variants [2, 3, 27, 46, 28, 49, 50] of ViViT have been introduced to improve the accuracy and reduce the computational cost. [51] presented a space-time mixing attention mechanism to reduce the complexity of the self-attention layers. TimeSFormer [46] introduced divided spatial and temporal attention to reduce the computational overhead. Then, it is further improved using the directed attention mechanism [28]. Then, [27] proposed a Multi-scale Vision Transformer (MViT) using multiscale feature hierarchies. Then, MViT-V2 [3] improves the performance of MViT by incorporating decomposed relative positional embeddings and residual pooling connections. Swin Video Transformer [2] has achieved state-of-the-art performance in action recognition by using shifted windows to limit the self-attention computational cost to local windows and also allow learning attention across windows. The recent SVFormer [52] has introduced a temporal warping augmentation to capture the complex temporal variation in videos for semi-supervised action recognition. Meanwhile, MTV [53] presented a Multiview Transformer to model different views of the videos with lateral connections to fuse information across views. Later, M&M Mix[54] further improved MTV by using multimodal inputs. TADA [55] proposed Temporally-Adaptive Convolutions (TAdaConv) to model complex temporal dynamics in videos. Inspired by the success of CLIP in vision-language pretraining [56, 57, 58], several studies have adopted this approach to video-language pretraining [5, 59, 60]. All-in-One [59] presented a unified approach to video-language pretraining by embedding raw video and textual inputs into joint representations with a unified network. EgoVLP [5] introduced Video-Language Pretraining to ego-centric video understanding. LF-VILA [60] presented a Multimodal Temporal Contrastive and Hierarchical Temporal Window Attention to model the long-form videos for video-language pretraining.
@@
Cross-view Video Learning The cross-view learning approaches have been exploited and proposed for several tasks, e.g., geo-localization [82, 83, 84, 85], semantic segmentation [86, 87, 88, 89, 90, 91]. Meanwhile, in video understanding tasks, several prior methods have alleviated the cross-view gap between exocentric and egocentric domains by using domain adaptation [16, 92, 93, 94], learning viewpoint-invariant [17, 95, 96, 97], or learning joint embedding [98, 99, 100, 101, 102]. Other works utilized generative models to synthesize the different viewpoints from a given image/video [103, 104, 105, 106]. However, these methods often require either a pair of data of both first and third views to learn the joint embedding or a share label domain when using domain adaptation [107, 108, 109].
@@ 4.2 Ablation Studies
Effectiveness of the scale $\alpha$ In this experiment, the metrics defined in Eqn. (12) and Eqn. (13) have been adopted to $\mathcal{D}_{x}$ and $\mathcal{D}_{a}$. The value $\alpha$ ranges from 0.0 to 2.0. When $\alpha=0.0$, it is equivalent to ViT simultaneously trained on both third-view and first-view datasets. As shown in Table 1, the mAP performance on the Charades-Ego benchmark is consistently improved when the value of $\alpha$ increases from $0.1$ to $0.75$ and achieves the best performance at the value of $\alpha=0.75$ and the mAP performance is $31.95\%$. Similarly, on the EPIC-Kitchen-55 benchmarks, the Top 1 and Top 5 accuracy is gradually improved w.r.t the increasing of $\alpha$ and reaches the maximum performance when the value of $\alpha$ is $1.50$ in which the Top 1 accuracy on EPIC Verb and EPIC Noun are $73.52\%$ and $68.19\%$. Then, the performance on both benchmarks steadily decreases when the value of $\alpha$ keeps increasing over the optimal point. Indeed, the variation in the video space is typically higher than in the attention maps due to the higher complexity of video data where the video data contains much more information, e.g., objects, humans, and interactions, etc.; meanwhile, the attention maps represent the focus of the models w.r.t model decisions. Thus, if the value of $\alpha$ is small, it could not represent the correct proportion of changes between videos and attention maps. Meanwhile, the higher value of $\alpha$ inclines to exaggerate the model focuses, i.e., attention maps, that results in the performance.
@@
Effectiveness of Transformer Layers This experiment studies the effectiveness of imposing the cross-view loss into attention maps of the Transformer layers. In this experiment, we adopt the optimal setting of the linear scale ($\alpha$) and correlation metrics ($\mathcal{D}_{x}$, $\mathcal{D}_{a}$) in the previous ablation studies. We consider four groups of Transformer layers, each consisting of three consecutive layers, i.e., Layer 1-3, Layer 4-6, Layer 7-9, and Layer 10-12. As experimental results in Table 3, the later Transformer layers of our model play an important role than the initial ones. In particular, when imposing the cross-view loss on only the first three Transformer layers, the performance of Charades-Ego has achieved 25.65% and the Top 1 accuracy of verb and noun predictions in EPIC-Kitchens-55 is 60.47% and 57.85%. Meanwhile, enforcing the cross-view self-attention loss into all attention layers brings better performance and achieves the best performance, i.e., the mAP of 31.95% on Charades-Ego and Top 1 accuracy of 73.52% and 68.19% on EPIC-Kitchens-55. Fig. 4 visualizes the attention maps of our model.
@@
Effectiveness of Paired and Unpaired Data. To illustrate the effectiveness of our proposed approach on paired data, we conduct an ablation study on the Charades Ego dataset. The Charades Ego provides a pair of both exocentric and egocentric videos. We evaluate our approach with two different settings of paired data, i.e., Charades-Exo $\to$ Charades-Ego, and unpaired data, i.e., Kinetics-400 $\to$ Charades-Ego. We conduct our experiments using different metrics of $\mathcal{D}_{x}$ and $\mathcal{D}_{a}$. As shown in Table 5, the performance of the model using paired data is slightly better than the one using unpaired data. However, since the data scale of Kinetics-400 is quite large compared to Charades-Exo, our approach using unpaired data achieves competitive performance. The experimental results have shown that our proposed approach remains efficient in both cases of paired and unpaired data. In addition, if the source data is relatively large, the model’s performance using unpaired data can even achieve competitive performance compared to the one using paired data.
Table 5: Effectiveness of Paired and Unpaired Data to the Charades-Ego Action Recognition Benchmarks.
$\mathcal{D}_{x}$ $\mathcal{D}_{a}$ Kinetics-400 Charades-Exo
$\ell_{2}$ $\mathcal{D}_{x}^{G}$ $\ell_{2}$ $\mathcal{D}_{a}^{JS}$to Charades-Ego $\to$ Charades-Ego
27.80 28.48
28.77 29.15
29.11 30.19
31.95 32.64
@@ 4.3 Comparisons with State-of-the-Art Results
Table 7: Comparisons on EPIC-Kitchen-55.
Method EPIC verbs EPIC nouns
Top 1 Top 5 Top 1 Top 5
ResNet-50 [15] 61.19 87.49 46.18 69.72
MViT-V2 [3] 55.17 89.87 56.59 79.40
Swin-B [2] 56.40 85.84 47.68 71.02
DANN [16] 61.27 87.49 45.93 68.73
Joint-Embed [17] 61.26 87.17 46.55 68.97
Ego-Exo + ResNet-50 [1] 62.83 87.63 48.15 70.28
Ego-Exo + SlowFast [1] 65.97 88.91 49.42 72.35
Ego-Exo* + ResNet-50 [1] 64.26 88.45 48.39 70.68
Ego-Exo* + SlowFast [1] 66.43 89.16 49.79 71.60
CVAR (Ours) 73.52 92.22 68.19 84.93